# People who frequently use ChatGPT for writing tasks are accurate and robust detectors of AI-generated text

### v1

@@ Abstract
In this paper, we study how well humans can detect text generated by commercial LLMs (GPT-4o, Claude-3.5-Sonnet, o1-Pro). We hire annotators to read 300 non-fiction English articles, label them as either human-written or AI-generated, and provide paragraph-length explanations for their decisions. Our experiments show that annotators who frequently use LLMs for writing tasks excel at detecting AI-generated text, even without any specialized training or feedback. In fact, the majority vote among five such “expert” annotators misclassifies only 1 of 300 articles, significantly outperforming most commercial and open-source detectors we evaluated even in the presence of evasion tactics like paraphrasing and humanization. Qualitative analysis of the experts’ free-form explanations shows that while they rely heavily on specific lexical clues (“AI vocabulary”), they also pick up on more complex phenomena within the text (e.g., formality, originality, clarity) that are challenging to assess for automatic detectors. We release our annotated dataset and code to spur future research into both human and automated detection of AI-generated text.
@@ 1 Introduction
Text generated by large language models (LLMs) is rampant in today’s world. [^1] For example, recent studies estimate that over 5% of recently-published Wikipedia articles Brooks et al. (2024) and 10% of PubMed abstracts published in 2024 Kobak et al. (2024) were written by AI. This state of affairs has spurred increased research into automatic detection of AI-generated text, which helps combat academic plagiarism Zhu et al. (2024) and fake content creation Gameiro et al. (2024). Unfortunately, automatic detectors suffer from low detection rates, poor robustness to evasion attempts, and limited explainability to end users Sadasivan et al. (2024); Ji et al. (2024). [^2] This paper focuses only on post-hoc detection of AI-generated text, which unlike watermarking Kirchenbauer et al. (2023) requires no cooperation from LLM providers.
@@
Unsurprisingly, annotators who rarely or never use LLMs are poor detectors of AI-generated text. However, we identify a subset of five high-performing annotators who frequently use LLMs for writing-related tasks. The majority vote of this subset of “expert” annotators fails to predict the correct label on only one out of 300 articles. We emphasize that we do not provide any training to our annotators: they are given no feedback on which articles they labeled incorrectly.
@@
The majority vote of our five expert annotators substantially outperforms almost every commercial and open-source detector we tested on these 300 articles, with only the commercial Pangram model Emi and Spero (2024) matching their near-perfect detection accuracy. In our most complex configuration, which requires detecting articles generated and humanized by o1-Pro, the expert majority vote is perfect (true positive rate of 100%), while open-source methods such as Binoculars (6.7%) and Fast-DetectGPT (23.3%) struggle. We conclude that hiring expert human annotators to perform detection is a viable strategy, particularly in high-stakes settings where explainability is critical.
@@
An analysis of our expert’s explanations reveals that usage of “AI vocabulary” (e.g., vibrant, crucial, significantly) form the most common giveaways. Close behind are formulaic sentence and document structures (e.g., optimistically vague introductions and conclusions) and originality (how creative or engaging an article is), while complex phenomena such as factuality and tone are also useful signals. We observe that neither paraphrasing nor humanization effectively removes all of these signatures; that said, these evasion tactics and defenses for them are still underexplored in the research community.
@@
We attempt to automatically replicate the decision-making process of our human experts by providing a candidate text to an LLM, along with a guidebook compiled from human expert explanations, and instructing it to use the guidebook to judge whether the text is written by an AI or by a human. This simple prompting strategy is competitive with existing detectors on easier configurations but struggles to detect humanized articles. Moving forward, we believe LLMs can be used in conjunction with high-performing detectors such as Pangram to offer an explanation of the judgment to end users.
Contributions:
Our study establishes the existence of an annotator population (frequent users of LLMs for writing-related tasks) that is highly capable of detecting AI-generated text without any additional training. The majority vote of five such annotators outperforms all automatic detectors we evaluated, even on paraphrased or humanized AI-generated articles. We release our annotated data and codebase to facilitate future research into both human and automatic detection of AI-generated text.
@@ 2 How good are humans at detecting AI-generated text?
Automatic detectors of AI-generated text suffer from low detection rates and poor robustness to simple perturbations of the text (e.g., paraphrasing), rendering them unreliable for real-world deployment Sadasivan et al. (2024). How well do human detectors stack up against their automatic counterparts? What features of a candidate text do humans find most predictive of AI writing? To answer these questions, we conducted five experiments that ask humans to judge increasingly more deceptive AI-generated texts (i.e., produced using more powerful base LLMs and evasion tactics).
@@
Each experiment consists of a batch of 60 articles, 30 of which are written by humans and the remaining 30 generated by an LLM. Each human-written article has a corresponding AI-generated counterpart with the same title and subtitle to control for high-level content and topics. The 60 articles are shown to annotators one at a time in a randomized order, and they are asked to provide the following for each article, depicted in Figure 1:
  1. 1.
    a binary label of whether the article was written by a human or an AI
  2. 2.
    a rating of their confidence in their choice on a scale from 1 (least confident) to 5 (most confident)
  3. 3.
    highlighted spans from the article that they used as clues to make their choice
  4. 4.
    a paragraph-length explanation of their choice
@@
We restrict our study to English nonfiction articles of fewer than 1K words geared towards lay audiences. [^3] Human-written articles in our experiments come from eight different American publications: Associated Press, Discover Magazine, National Geographic, New York Times, Reader’s Digest, Scientific American, Smithsonian Magazine, and Wall Street Journal. While we cannot make claims about human detection performance in other domains (e.g., scientific papers or social media posts), [^4] We do report results on §2.1 for fictional stories, which can be found in §C we focus on such articles because (1) malicious LLM users generate large volumes of fake articles, making it practically impactful; (2) such articles do not require specialized knowledge to understand or judge; and (3) many previous studies also ask humans to judge AI-generated text in this domain Ippolito et al. (2020); Brown et al. (2020); Clark et al. (2021); Puccetti et al. (2024).
@@
For each human-written article, we generate a corresponding AI article by prompting an LLM with the title, subtitle, desired length, and publication source of the original article. This ensures that AI-generated articles are directly comparable in topic and length, eliminating confounds from content variation. By constructing these human-AI article pairs with only authorship as the varying factor, we create minimal pairs that allow direct comparisons Gardner et al. (2020); Warstadt et al. (2020); Karpinska et al. (2022, 2024). We employ a within-subjects design for our experiments, where each annotator judges both the human-written and AI-generated articles. This reduces variability from individual differences and requires fewer annotators Allen (2017). To further minimize bias, annotators are unaware of the pairing, and the article order is randomized. An example article generation prompt is below:
Title: The Science That Makes Baseball Mud ‘Magical’
Article:
For experiments 3 and 5, we modify the prompt to include perturbations designed to evade detectors (paraphrasing and humanization). Comparisons of article lengths between human and AI-generated articles for each experiment are reported in Table 1 and further detailed in §B.1.555We set temperature=0 for Experiments 1-3 and use ChatGPT (Pro) interface for Experiments 4-5 as the API for o1 was not yet available.
Exp # Model Human-written (n=30) AI-generated (n=30)
1 GPT-4o 652.5170.9 625.1126.9
2 Claude-3.5-Sonnet 793.1170.3 710.8133.5
3 GPT-4o paraphrased 777.8175.9 654.9110.6
4 o1-Pro 728.4155.1 813.8233.6
5 o1-Pro humanized 739.0128.3 815.6171.2
@@
We recruited annotators via Upwork, [^6] https://www.upwork.com/ all of whom identified themselves as native English speakers. All annotators were required to read the project guidelines and sign a consent form prior to the task, provided in §A. Additionally, we surveyed each annotator to determine their level of education, profession, English dialect, and familiarity with LLMs, including how they use these models (see Table 5 for details). They were compensated \$2 per candidate text, resulting in an hourly rate of about \$15 - \$30 USD. Ultimately, we collect 1740 annotations from 9 annotators at a total cost of \$4.9K USD.
Evaluation metrics:
We evaluate both human and automatic detectors on True Positive Rate (TPR), which measures the percentage of AI-generated articles that are successfully detected, as well as False Positive Rate (FPR), which measures the percentage of human-written articles incorrectly detected as AI-generated. Prior work in automatic detection reports either TPR at a fixed low FPR such as 1% Krishna et al. (2023); Hans et al. (2024); Dugan et al. (2024) or AUROC Mitchell et al. (2023); Hu et al. (2023); Bao et al. (2023), which we cannot easily do with human annotators. For fair comparison, we report TPR and FPR of automatic detectors using recommended thresholds whenever possible, otherwise using a threshold calibrated to an FPR of 5% on a held-out set of 40 human-written articles.
@@
Is there a population of humans that is highly accurate at our detection task, and if so, are there any commonalities between them? To explore these questions, we first ask five annotators with varying backgrounds and levels of expertise with LLMs to attempt a batch of 60 articles (half generated by GPT-4o 2024-08-06 and half human-written as described above). We observe both ends of the spectrum in terms of performance: annotators who rarely or never use LLMs for writing tasks performed near random, while one annotator who uses LLMs (and edits LLM-generated text) daily performed almost perfectly. More details on our initial study can be found in §C. Inspired by this result, we recruit four more annotators with similar backgrounds as our top performer, all of whom perform similarly well on the same batch of articles.
Annotators with limited LLM experience are unreliable detectors:
The four annotators who self-report either not using LLMs at all, or using LLMs only for non-writing tasks, detect AI-generated text at a similar rate to random chance, achieving an average TPR of 56.7% and FPR of 52.5% (Table 2). [^7] We report the average TPR and FPR since there were only four non-expert annotators. For the remainder of the paper, we will report the majority vote of the five expert annotators (i.e., at least three out of five agreed). While performance among these annotators varies greatly, no single annotator could reliably detect AI-generated text. We term this population nonexperts for the rest of this paper.
In contrast, the five annotators who have significant experience with using LLMs for writing-related tasks are able to detect AI-generated text very reliably, achieving a TPR of 92.7%. The average FPR for this population of annotators was 3.3%, meaning that they rarely mistake human-written text as AI-generated. The majority vote out of these five annotators correctly determined authorship of all 60 articles; the “GPT-4o” column of Table 3 contains more details. We term this population experts for the rest of this paper; all subsequent experiments rely on only these five expert annotators.
What do expert annotators see that nonexperts don’t?
To understand why experts far outperformed nonexperts at detecting AI-generated text, we analyze the comments each annotator provided in their explanations. Overall, nonexperts often mistakenly fixate on certain linguistic properties compared to experts. One example is vocabulary choice, where nonexperts take the inclusion of any “fancy” or otherwise low-frequency word types as signs of AI-generated text; in contrast, experts are much more familiar with exact words and phrases overused by AI (e.g., testament, crucial). [^8] A complete list of “AI vocab” found in the detection guide (Table 11) is detailed in Table 12. Nonexperts also believe that human authors are more likely to form grammatically-correct sentences and thus attribute run-on sentences to AI, while experts realize the opposite is true: humans are more likely than AI to use ungrammatical or run-on sentences. Finally, nonexperts attribute any text written in a neutral tone to AI, which results in many false positives because formal human writing is also often neutral in tone.
@@
Despite the change in model, experts are reliable at detecting AI-generated content. TPR is almost unchanged from Experiment 1 and the expert majority vote is again 100% accurate (“Claude-3.5-Sonnet” column of Table 3), with two annotators achieving a perfect score individually. We note that one expert (Annotator 2) achieved higher performance in the experiments where an OpenAI model was used, with their TPR falling to 80% in this experiment. Annotator 2 did not have any prior experience using Claude, although three other annotators who were also unfamiliar with Claude achieved TPRs of at least 96.7%. However, Annotator 2 focuses more heavily on the existence of “AI vocab” in the article than the other annotators, and they were not familiar with the distribution of words that are overused by Claude.
@@
Users of LLMs who hope to evade detection often resort to simple post-hoc modifications of LLM outputs such as paraphrasing. These tactics significantly drop detection rates of automatic methods Krishna et al. (2023); Sadasivan et al. (2024), but how do they affect our human experts? We evaluate the five expert annotators on a third set of 60 articles, where the 30 AI-generated articles go through an additional round of paraphrasing, and we show that they maintain high TPR and low FPR.
Creating paraphrased AI-generated articles:
We follow the sentence-by-sentence paraphrasing approach with LLMs outlined by Chang et al. (2024) to create a more difficult set of AI articles for this experiment. Additionally, since our experts easily identified AI-generated named entities (e.g., Sarah Thompson) in the previous experiments, we extract named entities from the human-written article and instruct GPT-4o to include these names in its generated article. We use GPT-4o as both the article generation model and the paraphrasing model. See Table 10 for paraphrasing prompts used in this experiment.
@@
Overall, TPR and FPR remain almost unchanged in this experiment compared to the first two, suggesting that paraphrasing is not an effective attack on expert human detectors (see “gpt-4o paraphrased” column in  Table 3). The majority vote among the five experts again correctly detects the labels of all 60 articles. Annotators continue to pick up on many of the same clues within the paraphrased articles that were also apparent in Experiments 1 & 2, such as high frequency of “AI vocab” (even after paraphrasing), formulaic sentence structures, and cheerful summary conclusions.
@@
While we were in the middle of conducting experiments, OpenAI released their o1 model OpenAI (2024), ushering in a new paradigm of test-time scaling. This release offered us a unique opportunity: since none of our experts had been previously exposed to a model with such advanced reasoning capabilities Zhong et al. (2024), how well would they do at detecting its output? Interestingly, our experts remain reliable detectors of articles generated by o1-Pro (using the same prompt as in Experiments 1 and 2),999o1-Pro is available through the ChatGPT Pro subscription. OpenAI states o1-Pro produces more reliably accurate and comprehensive responses than o1. although their comments show that they often perform a more detailed analysis of articles to make their judgments.
@@
For four of five experts, detection ability of o1-generated content remains largely consistent with prior experiments, with a majority vote TPR of 96.7% and FPR of 0% (see “o1-Pro” column of Table 3). [^10] One AI-generated article was incorrectly marked by the expert majority vote as human-written. Analyzing their explanations reveals that the three annotators who rely most heavily on vocabulary and sentence structure did not find enough clues in those categories to judge the article as AI-generated. The full text of the article is in Table 14. Interestingly, Annotator 3’s TPR drops considerably, as they prioritize signs of human writing over signs of AI writing when making their judgments. Average confidence dropped to 4.21 out of 5, compared to average confidence of 4.39, 4.38, and 4.48 from Experiments 1,2, & 3 respectively (see Figure 2 for details). The drop in confidence, as well as verbal feedback from our annotators, demonstrates the increased challenge posed by o1-generated articles. However, their steady aggregate performance highlights that even new model paradigms are still detectable by human experts.
@@
As the deployment of automatic AI detectors increases, many users are attempting to evade them by using humanization methods, which explicitly attempt to make AI-generated text more human-like Wang et al. (2024a, b) unlike more generic attacks such as paraphrasing. Are such methods effective at evading our human experts? Since no well-established humanization methods exist, we first create our own humanizer by prompting o1-Pro with a detailed set of instructions derived from expert comments written for Experiments 1-4. Despite considerably degrading performance of many automatic detectors, the majority of our experts remain robust to humanization (“o1-Pro humanized” column of Table 3).
@@
We modify the article generation prompt used in prior experiments to include detailed instructions on specific AI signatures to avoid when generating the article. To obtain these instructions, we consult our experts directly, paying them \$45 each to provide us with a list of clues that they look for during detection. We then manually organize these disparate clues into a unified “guidebook” with different sections (e.g., vocabulary, grammar, tone, introductions, conclusions), where each section provides explanations and examples of how AI writing differs from human writing (see Table 11 for a truncated version of the guidebook). Then, we prompt o1-Pro with the guidebook and an instruction to generate an article that would evade a detector that was following the guidebook (see Table 13 for prompt). [^11] We only use o1-Pro for this experiment as the previous experiments show it was more difficult to detect than GPT-4o and Claude-3.5-Sonnet. Through this process, we create a new batch of 60 articles, 30 of which are generated by our humanization prompt using o1-Pro. More details on the development of our humanization method can be found in §B.2.
@@
Despite our best efforts to generate articles that our experts would find undetectable, most of their detection rates remain largely unchanged from prior experiments, and the expert majority vote is again perfect on all 60 articles. Annotator 3, who struggled to detect non-humanized o1-Pro articles in Experiment 4, performed remarkably poorly on this batch, achieving a TPR of zero by marking almost every article as human-written. While detection rates remained steady, overall confidence dropped, with 15.1% of annotations having a confidence of 1, showing the increased difficulty presented by the humanized articles (see Figure 2).
@@
Although our experts reported this round to be the most challenging, they remained consistent in the types of clues used from prior rounds, demonstrating that the humanization prompt did not remove the entirety of the “AI signature” from the articles. For example, while humanization increases variety in AI-generated names, often even including people relevant to the article who exist in real life, our experts found that humanized articles use titles for these people (e.g., Dr., Prof.) much more frequently than in human-written articles. Overall, our results highlight the need for increased public research on humanization, as more advanced methods can challenge both human and automated detectors while offering better training data for more robust detection systems.
3 Fine-grained analysis of expert performance

### v2

@@ Abstract
In this paper, we study how well humans can detect text generated by commercial LLMs (GPT-4o, Claude-3.5-Sonnet, o1-Pro). We hire annotators to read 300 non-fiction English articles, label them as either human-written or AI-generated, and provide paragraph-length explanations for their decisions. Our experiments show that annotators who frequently use LLMs for writing tasks excel at detecting AI-generated text, even without any specialized training or feedback. In fact, the majority vote among five such “expert” annotators misclassifies only 1 of 300 articles, significantly outperforming most commercial and open-source detectors we evaluated even in the presence of evasion tactics like paraphrasing and humanization. Qualitative analysis of the experts’ free-form explanations shows that while they rely heavily on specific lexical clues, they also pick up on more complex phenomena within the text that are challenging to assess for automatic detectors. We release our annotated dataset and code to spur future research into both human and automated detection of AI-generated text.
@@ 1 Introduction
AI-generated text is rampant in today’s world, spurring research into automatic detectors to help combat academic plagiarism Zhu et al. (2024) and fake content creation Gameiro et al. (2024). Unfortunately, automatic detectors suffer from low detection rates, poor robustness to evasion attempts, and limited explainability to end users Sadasivan et al. (2024); Ji et al. (2024). [^1] This paper focuses only on post-hoc detection of AI-generated text, which, unlike watermarking Kirchenbauer et al. (2023) requires no cooperation from LLM providers.
@@
Unsurprisingly, annotators who rarely or never use LLMs are poor detectors of AI-generated text; in fact, they overestimate their own ability to perform the task by providing high confidence scores for their decisions. We identify a subset of five high-performing annotators who frequently use LLMs for writing tasks (e.g., editing, copywriting, creative writing). The majority vote of this subset of “expert” annotators fails to predict the correct label on only one out of 300 articles. We emphasize that we do not provide any training to our annotators: they are given no feedback on which articles they labeled incorrectly.
@@
The majority vote of our five expert annotators outperforms almost every commercial and open-source detector we tested on these 300 articles, with only the commercial Pangram model Emi and Spero (2024) matching their near-perfect detection accuracy. In our most complex configuration, which requires detecting articles generated and humanized by o1-Pro, the expert majority vote is perfect (true positive rate of 100%), while open-source methods such as Binoculars (6.7%) and Fast-DetectGPT (23.3%) struggle. We conclude that hiring expert human annotators to perform detection is a viable strategy, particularly in high-stakes settings where explainability is critical.
@@
An analysis of our expert’s explanations reveals that usage of “AI vocabulary” (e.g., vibrant, crucial, significantly) form the most common giveaways. Close behind are formulaic sentence and document structures (e.g., optimistically vague conclusions) and originality (how creative or engaging an article is). We observe that neither paraphrasing nor humanization effectively removes all of these signatures; that said, these evasion tactics and defenses for them are still underexplored in the research community.
@@
We attempt to automatically replicate the decision-making process of our human experts by providing a candidate text to an LLM, along with a guidebook compiled from human expert explanations, and instructing the model to use the guidebook to judge whether the text is written by an AI or by a human. This simple prompting strategy is competitive with existing detectors on easier configurations but struggles to detect humanized articles. Moving forward, we believe LLMs can be used in conjunction with high-performing detectors such as Pangram to offer an explanation of the judgment to end users.
@@ 2 How good are humans at detecting AI-generated text?
How well do human detectors stack up against their automatic counterparts? What features do humans find most predictive of AI writing? To answer these questions, we conducted five experiments that ask humans to judge increasingly more deceptive AI-generated texts.
@@
Each experiment consists of 60 articles, 30 human-written and 30 LLM-generated. Each human-written article has a corresponding AI-generated counterpart with the same title and subtitle to control for high-level content and topics. Articles are shown to annotators in a randomized order. For each, they provide (1) a binary label (human-written or AI-generated), (2) a confidence rating on a scale from 1 (least confident) to 5 (most confident), (3) highlighted spans used as clues, and (4) a paragraph-length explanation. We evaluate with True Positive Rate (TPR), which measures the percentage of AI-generated articles that are successfully detected, as well as False Positive Rate (FPR), which measures the percentage of human-written articles marked as AI-generated. [^2] Prior work in automatic detection reports either TPR at a fixed low FPR such as 1% Krishna et al. (2023); Hans et al. (2024); Dugan et al. (2024) or AUROC Mitchell et al. (2023); Hu et al. (2023); Bao et al. (2023), which we cannot easily do with human annotators.
@@
We restrict our study to English nonfiction articles of fewer than 1K words. [^3] Human-written articles in our experiments come from eight different American publications: Associated Press, Discover Magazine, National Geographic, New York Times, Reader’s Digest, Scientific American, Smithsonian Magazine, and Wall Street Journal. While we cannot make claims about human detection performance in other domains (e.g., scientific papers or social media posts), [^4] We report results on §2.1 for fictional stories in §C. we focus on such articles because (1) malicious LLM users generate large volumes of fake articles, making it practically impactful; (2) such articles do not require specialized knowledge to understand or judge; and (3) many previous studies also ask humans to judge AI-generated text in this domain Ippolito et al. (2020); Clark et al. (2021); Puccetti et al. (2024).
@@
For each human-written article, we generate a corresponding AI article by prompting an LLM with its title, subtitle, desired length, and publication source. This ensures that AI-generated articles are directly comparable in topic and length, [^5] We include a word count target to ensure comparable lengths between article pairs, reported in Table 6. By constructing these human-AI article pairs with only authorship as the varying factor, we create minimal pairs that allow direct comparisons Gardner et al. (2020); Warstadt et al. (2020); Karpinska et al. (2022, 2024). For experiments 3 and 5, we modify the prompt to include perturbations designed to evade detectors (paraphrasing and humanization). We employ a within-subjects design for our experiments, where each annotator judges both the human-written and AI-generated articles. This reduces variability from individual differences and requires fewer annotators Allen (2017). We further minimize bias by randomizing article order and ensuring annotators are unaware of the pairing. [^6] We set temperature=0 for Experiments 1-3 and use ChatGPT (Pro) interface for Experiments 4-5 as the API for o1 was not yet available.
@@
We recruited annotators via the Upwork freelancing platform, all of whom self-identified as native English speakers. [^7] See Table 4 for survey details. All annotators were required to read the project guidelines and sign a consent form prior to the task, provided in §A. Additionally, we surveyed each annotator to determine their level of education, profession, English dialect, and familiarity with LLMs, including how they use these models (see Table 4 for details). They were compensated \$2 per candidate text, resulting in an hourly rate of about \$15 - \$30 USD. Ultimately, we collect 1740 annotations from 9 annotators at a total cost of \$4.9K USD.
Justification of sample size:
Most experiments in this paper are conducted by five expert annotators (Experiment 1 includes four nonexpert annotators). [^8] See §A.1 for details on expert recruitment. While a larger sample size would strengthen population-level claims; prohibitive costs [^9] It cost us \$865 USD per annotator for five experiments. limit us to the current sample size. We employ a within-subjects design for our experiments, where each annotator judges both the human-written and AI-generated articles, reducing individual variability and requiring fewer annotators Allen (2017). Our set of five expert annotators provide an insightful view of the human upper bound at AI-generated text detection without additional training. Furthermore, collecting detailed explanations of these annotators’ decision-making process, while expensive, enables fine-grained qualitative analysis and ensures that annotators actually read the texts. [^10] Collecting explanations mitigates issues raised by  Karpinska et al. (2021), whose Mechanical Turk annotators “read” long-form texts in a matter of mere seconds before evaluation.
@@
Is there a population of humans that is highly accurate at our detection task, and if so, are there any commonalities between them?
We first ask five annotators with varying backgrounds and levels of expertise with LLMs to attempt a batch of 60 articles. We observe both ends of the spectrum in terms of performance: annotators who rarely or never use LLMs for writing tasks performed near random, while one annotator who uses LLMs daily to edit LLM-generated text performed almost perfectly. More details on our initial study can be found in §C. Inspired by this result, we recruit four more annotators with similar backgrounds as our top performer, who all of whom perform similarly well on the same batch of articles.
@@
Annotators with limited LLM experience are unreliable detectors:
The four annotators who self-report either not using LLMs at all, or using LLMs only for non-writing tasks, detect AI-generated text at a similar rate to random chance, achieving an average TPR of 56.7% and FPR of 52.5% (Table 1). [^11] We report average TPR and FPR instead of majority vote since there were only four nonexpert annotators. For the remainder of the paper, we report the majority vote of the five expert annotators (i.e., at least three out of five agreed). Despite their poor performance, these annotators reported high confidence (avg. 4.03) in their decisions, suggesting they may be overestimating their detection abilities. We term this population nonexperts for the rest of this paper.
In contrast, the five annotators who have significant experience with using LLMs for writing-related tasks are able to detect AI-generated text very reliably, achieving a TPR of 92.7%. [^12] We compare expert vs. nonexpert clues in §D.2. The average FPR for this population of annotators was 3.3%, meaning that they rarely mistake human-written text as AI-generated. The majority vote out of these five annotators correctly determined authorship of all 60 articles; the “GPT-4o” column of Table 2 contains more details. We term this population experts for the rest of this paper; all subsequent experiments rely on only these five expert annotators.
What do expert annotators see that nonexperts don’t?
To understand why experts far outperformed nonexperts at detecting AI-generated text, we analyze the comments each annotator provided in their explanations. Overall, nonexperts often mistakenly fixate on certain linguistic properties compared to experts. One example is vocabulary choice, where nonexperts take the inclusion of any “fancy” or otherwise low-frequency word types as signs of AI-generated text; in contrast, experts are much more familiar with exact words and phrases overused by AI (e.g., testament, crucial). [^13] A complete list of “AI vocab” found in the detection guide (Table 11) is detailed in Table 12. Nonexperts also believe that human authors are more likely than AI to form grammatically-correct sentences, while experts realize the opposite is true: humans actually make more grammatical errors. Finally, nonexperts attribute any text written in a neutral tone to AI, resulting in many false positives.
@@
Despite the change in model, experts are reliable at detecting AI-generated content. TPR is almost unchanged from Experiment 1 and the expert majority vote is again 100% accurate (“Claude-3.5-Sonnet” column of Table 2), with two annotators achieving a perfect score individually; however, we find one annotator performs worse on Claude-generated texts than those of GPT-4o. [^14] This annotator’s performance is detailed in §D.3. We note that one expert (Annotator 2) achieved higher performance in the experiments where an OpenAI model was used, with their TPR falling to 80% in this experiment. Annotator 2 did not have any prior experience using Claude, although three other annotators who were also unfamiliar with Claude achieved TPRs of at least 96.7%. However, Annotator 2 focuses more heavily on the existence of “AI vocab” in the article than the other annotators, and they were not familiar with the distribution of words that are overused by Claude.
@@
Users of LLMs who hope to evade detection often resort to simple post-hoc modifications of LLM outputs such as paraphrasing. These tactics significantly drop detection rates of automatic methods Krishna et al. (2023); Sadasivan et al. (2024), but how do they affect our human experts? We evaluate the five expert annotators on a third set of 60 articles, where the 30 AI-generated articles go through an additional round of paraphrasing, [^15] Details on paraphrasing setup are in §B.2. showing they maintain high TPR and low FPR.
@@
Overall, TPR and FPR remain almost unchanged in this experiment compared to the first two, suggesting that paraphrasing is not an effective attack on expert human detectors (see “gpt-4o paraphrased” column in Table 2). The majority vote among the five experts again correctly detects the labels of all 60 articles.
@@
While we were in the middle of conducting experiments, OpenAI released their o1 model OpenAI (2024), ushering in a new paradigm of test-time scaling. This release offered us a unique opportunity: since none of our experts had been previously exposed to a model with such advanced reasoning capabilities Zhong et al. (2024), how well would they do at detecting its output? Interestingly, our experts remain reliable detectors of articles generated by o1-Pro (using the same prompt as in Experiments 1 and 2),161616o1-Pro is available through the ChatGPT Pro subscription. OpenAI states o1-Pro produces more reliably accurate and comprehensive responses than o1. although their comments show that they often perform a more detailed analysis of articles to make their judgments.
@@
For four of five experts, detection ability of o1-generated content remains largely consistent with prior experiments, with a majority vote TPR of 96.7% and FPR of 0% (see “o1-Pro” column of Table 2). [^17] One AI-generated article was misclassified by the expert majority vote. Analyzing their explanations, the three annotators most focused on vocabulary and sentence structure did not find enough clues in those categories to judge the article as AI-generated. The full text of the article is in Table 14. Average confidence dropped to 4.21 out of 5, compared to average confidence of 4.39, 4.38, and 4.48 from Experiments 1,2, & 3 respectively (see Figure 2 for details). Verbal feedback from experts, along with the drop in confidence, demonstrates the increased challenge posed by o1-generated articles. However, their steady aggregate performance highlights that even new model paradigms are still detectable by human experts.
@@
Many users attempt to evade detectors via humanization methods, which explicitly attempt to make AI-generated text more human-like Wang et al. (2024a, b). Are such methods effective at evading our human experts? Since no well-established humanization methods exist, we first create our own humanizer by prompting o1-Pro with a detailed set of instructions derived from expert explanations from Experiments 1-4. Despite considerably degrading performance of many automatic detectors, humanization does not fool our experts in aggregate (“o1-Pro humanized” column of Table 2).
@@
We modify the article generation prompt used in prior experiments to include instructions on specific AI signatures to avoid. To obtain these instructions, we pay our experts \$45 each to provide us with a list of clues that they look for during detection. We then then manually organize these disparate clues into a “guidebook” with different sections (e.g., vocabulary, tone) that each contain explanations and examples of how AI writing differs from human writing (see Table 11 for a truncated version of the guidebook). Then, we prompt o1-Pro with the guidebook and an instruction to generate an article that would evade a detector that was following the guidebook (see Table 13 for prompt). [^18] We only use o1-Pro for this experiment as the previous experiments show it was more difficult to detect than GPT-4o and Claude-3.5-Sonnet. More details on the development of our humanization method can be found in §B.2.
@@
Despite our best efforts to generate articles that our experts would find undetectable, most of their detection rates remain largely unchanged from prior experiments, and the expert majority vote is again perfect on all 60 articles. Annotator 3, who struggled to detect non-humanized o1-Pro articles in Experiment 4, performed remarkably poorly on this batch, achieving a TPR of zero by marking almost every article as human-written. While detection rates remained steady, overall confidence dropped, with 15.1% of annotations having a confidence of 1, showing the increased difficulty presented by the humanized articles (Figure 2).
@@
The clues used by experts remained consistent from prior experiments, showing that humanization does not completely remove the “AI signature” from texts. For example, while humanization increases variety in AI-generated names, often even including people relevant to the article who exist in real life, our experts found that humanized articles use titles for these people (e.g., Dr., Prof.) much more frequently than in human-written articles. Our results highlight the need for increased public research on humanization, as more advanced methods can challenge both human and automated detectors while offering better training data for more robust detection systems.
3 Analysis of expert performance
We first provide in-depth comparisons between experts and automatic detectors. [^19] We release our dataset of annotations under under the MIT License. Next, we perform a fine-grained coding of expert explanations into different categories (e.g., vocabulary, originality, formality), which allows us to examine what they focus on during the different experiments. Finally, we analyze differences between annotators, highlighting implications for the future training of human annotators for AI-generated text detection.